OpenStudy (anonymous):

can someone help me figure out x. 3267 = 27(10)^x

sam (.sam.):

3267=27(10)^(x) Since x is on the right-hand side of the equation, switch the sides so it is on the left-hand side of the equation. 27(10)^(x)=3267 Divide each term in the equation by 27. (27(10)^(x))/(27)=(3267)/(27) Simplify the left-hand side of the equation by canceling the common factors. (10)^(x)=(3267)/(27) Simplify the right-hand side of the equation by simplifying each term. (10)^(x)=121 \[\log_{10}(121)=x \] x=2.0829

OpenStudy (anonymous):

Thanks a lot i appreciate it.